A rocket is fired with an initial velocity of 100m/s at an angle of 55° above the horizontal it explode on the mountain side 12second after its fired what is the x&y coordinate of the rocket relative to its firing point

1
Expert's answer
2022-05-18T08:31:50-0400
"x=v_0\cos\theta t=100*\cos55^\circ*12=688\:{\rm m}\\
y=v_0\sin\theta t-gt^2/2\\
=100*\sin55^\circ*12-9.8*12^2/2=277\:\rm m"